OpenStudy (aem9296):

Solve using logarithms 7^(x-2)=5^(3-x)

OpenStudy (babalooo):

when you log something like say \[x^{5}\] and you log it to make \[\log x^{5} \] you can pull the exponent down to the front like so \[5\log x \]

OpenStudy (babalooo):

if that is the case, you can pull the x terms in the exponents and solve.

OpenStudy (babalooo):

pull the x terms in the exponents to the front*

OpenStudy (aem9296):

but doesn't that kinda take you backwards since you have to solve for x?

OpenStudy (babalooo):

but you can't exactly solve for x while having them in the exponents can you? if you pull the (x-2)to the front (x-2)log7 = (3-x)log5, you can isolate like terms that contain x and the rest would just be in decimal numbers

OpenStudy (aum):

\(\large 7^{x-2} = 5^{3-x}\) Take logs on both sides: \(\large \log(7^{x-2}) = \log(5^{3-x})\) \(\large (x-2)\log(7) = (3-x)\log(5)\) \(\large x\log(7)-2\log(7) = 3\log(5)-x\log(5)\) Add \(\large x\log(5)\) to both sides: \(\large x\log(7) + x\log(5) -2\log(7) = 3\log(5)\) Add \(\large 2\log(7)\) to both sides: \(\large x\log(7) + x\log(5) = 3\log(5) + 2\log(7) \) \(\large x(\log(7) + \log(5)) = 3\log(5) + 2\log(7) \) \(\large x\log(35) = 3\log(5) + 2\log(7) \) \(\large x = \Large \frac{3\log(5) + 2\log(7)}{\log(35)} \)